OpenStudy (coolguy34):

I need help. Solve for the variable: 18b ≤ -3

Mehek (mehek14):

to get rid of 18 from b, divide 18 \(\dfrac{18}{18}b≤\dfrac{-3}{18}=b≤\dfrac{-3}{18}\)

OpenStudy (coolguy34):

please put that is simpler terms that way i can copy on to my document

Mehek (mehek14):

\(b≤\dfrac{-3}{18}\) ?

OpenStudy (coolguy34):

thanks